OpenStudy (anonymous):

Guass Jordan elimination. x1 + 6x2 - x3 - 4x4 = 0 -2x1 - 12x2 + 5x3 + 17x4 = 0 3x1 + 18x2 - x3 - 6x4 = 0

OpenStudy (anonymous):

first set it up as a matrix can you do that?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

1 6 -1 -4 0 -2 -12 5 17 0 3 18 -1 -6 0

OpenStudy (anonymous):

i got 480

OpenStudy (anonymous):

jokes

OpenStudy (anonymous):

okay so this is a 3x4 matrix so you need to put it in ref so start eliminating the bottom first

OpenStudy (anonymous):

I would do these operations first \[R_2 \rightarrow 2R_1 + R_2\] \[R_3 \rightarrow -3R_1 + R_3\]

OpenStudy (anonymous):

ok let me show that step..

OpenStudy (anonymous):

1 6 -1 -4 0 0 0 3 9 0 0 0 2 6 0 Then 1/3R2 = R2 and 1/2 R3 = R3

OpenStudy (anonymous):

yup thats right

OpenStudy (anonymous):

ok I have the next step I think...

OpenStudy (anonymous):

1 6 -1 -4 0 0 0 1 3 0 0 0 1 3 0 now R3 + (-1)R2 = R3

OpenStudy (anonymous):

oops I left something out sorry...let me fix that.

OpenStudy (anonymous):

yea i was about to point out an error in that matrix you have

OpenStudy (anonymous):

and yea that operation you metioned is correct

OpenStudy (anonymous):

Now Im confused..where is the error?

OpenStudy (anonymous):

wait nvm ignore that my mistake keep going

OpenStudy (anonymous):

First R1 + R3 = R1

OpenStudy (anonymous):

to get...

OpenStudy (anonymous):

1 6 0 -1 0 0 0 1 3 0 0 0 1 3 0

OpenStudy (anonymous):

then R3 + (-1)R2 = R3

OpenStudy (anonymous):

good so far

OpenStudy (anonymous):

to get..

OpenStudy (anonymous):

1 6 0 -1 0 0 0 1 3 0 0 0 0 0 0

OpenStudy (anonymous):

and this is the part I really need help with......past this.

OpenStudy (anonymous):

okay technically your done reducing alrready now all you need to do is put it in terms of x1, x2 etc

OpenStudy (anonymous):

you get what im saying i do the first row for you x1 + 6x2 - x3 = 0

OpenStudy (anonymous):

Im solving just a sec...

OpenStudy (anonymous):

-x4 sorry

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

so how does.... x1 = -6R + S x2 = R x3 = -3S x4 = S ......look?

OpenStudy (anonymous):

Im not sure if Im doing this part correctly...

OpenStudy (anonymous):

how you get x3 = -3s?

OpenStudy (anonymous):

lets do this step by step

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

lol nvm your right sorry

OpenStudy (anonymous):

did the question also asked to put it in parameterized vector form?

OpenStudy (anonymous):

no. What does that mean?

OpenStudy (anonymous):

so all you do to finish it up is put it like this: \[\left(\begin{matrix}x_1 \\ x_2 \\ x_3 \\ x_4\end{matrix}\right) = \left(\begin{matrix}-6R + S\\ R \\ -3S \\ S\end{matrix}\right) = \left(\begin{matrix}-6R \\ R \\ 0 \\ 0\end{matrix}\right) + \left(\begin{matrix}0 \\ 0 \\ -3S \\ S\end{matrix}\right)\]

OpenStudy (anonymous):

= \[\left(\begin{matrix}-6 \\ 1 \\ 0 \\ 0\end{matrix}\right)R + \left(\begin{matrix}0 \\ 0 \\ -3 \\ 1\end{matrix}\right)S \] and your done

OpenStudy (anonymous):

aw dam i made a mistake its supposed to be a 1 instead of a 0 on the first row for the vector with the S

OpenStudy (anonymous):

I think I still follow it though. Thank you so much for walking me through this one, this is a new concept for me, seeing lineear algebra for the first time.

OpenStudy (anonymous):

no prob glad to help its all about practice and you'll get used to it